OpenStudy (anonymous):

fine the difference quotient for the function f(x)=-8x+4 simplify it ?

OpenStudy (unklerhaukus):

\[f(x)=-8x+4\] \[f(x+h)=-8(x+h)+4\] \[\frac{f(x+h)-f(x)}{h}=\]

OpenStudy (unklerhaukus):

What do you get?

OpenStudy (anonymous):

\[fx+fh-fx divh\]

OpenStudy (anonymous):

i meant fx+fh-fx divided by h

OpenStudy (unklerhaukus):

no f(x+h) is not equal to f(x) + f(h)

OpenStudy (unklerhaukus):

\[\frac{f(x+h)-f(x)}{h}=\frac{[-8(x+h)+4]-[-8x+4]}{h}= \cdots\]

OpenStudy (unklerhaukus):

Simplify!